How can three resistors of resistances 2 Ω , 3 Ω , and 6 Ω be connected to give a total resistance of 1 Ω?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


When R1,R2, R3 are connected in parallel we get
1/R = 1/R1 + 1/R2 + 1/R3
= 1/2 + 1/3 + 1/6
= 1 ohm.

Thanks